OpenStudy (anonymous):

What is the surface area of this triangular prism with a length of 10 cm, a base of 6 and a height is 4.

OpenStudy (anonymous):

Please help me!!! I've been stuck on this for a while now, I just don't get the whole surface area thing😢🌽😭

Directrix (directrix):

The surface area is the sum of the areas of the three lateral faces added to each of the bases to get the total area. Total area = p*h + 2*B where p is the perimeter of the triangular base, h is the height of the prism, and B is the area of the Base. So ...

Directrix (directrix):

I just realized that I need to see the diagram of the prism. Will you copy it into this thread? @jarchinard

OpenStudy (anonymous):

Thnx so much!!!!!😃😃

Directrix (directrix):

Will you copy the diagram into this thread? @jarchinard

OpenStudy (anonymous):

Yeah sure!!!

OpenStudy (anonymous):

Sure! Here it is;

Directrix (directrix):

Total area = p*h + 2*B

OpenStudy (anonymous):

triangles + base + sides \[2*3*4 + 6*10 + 2*5*10 = 184\]

OpenStudy (anonymous):

I'm sorry this isn't a triangular prism, I meant rectangular